# Kandi, Sangareddy, Telengana 502285, India

### v1

@@ 4 Volume complexity
$\mathcal{C}_{V}(t)=2\int_{r_{\text{min}}}^{\infty}dr\int_{0}^{rx_{1}}dz\frac{\ell^{2}}{(\ell+z)^{3}}f(r,z).$ (48)
@@
We can proceed with the extremization of this volume functional. For the extermination, we can think of the volume functional as a Lagrangian $\mathcal{L}(\dot{t},\dot{r},r)$ which is in general a function of a function $r(\lambda)$ and $t(\lambda)$ but as in this case it is independent of $t$ we will have a constant of motion,
@@
where, $H_{0}(r)=\frac{r^{2}}{l_{3}^{2}}+\kappa$. If we consider $\kappa=-1$, we get the analog of the classical BTZ black hole result Carmi:2017jqz with the classical horizon at $r_{+}=\ell_{3}$ hidding the charged defect. Below, we focus on the quantum correction and late-time growth of complexity in this case. For the $\kappa=1$ case, one can proceed similarly, with the only difference coming from the fact that there is no horizon classically i.e. there is no real root of $H(r)$. The horizon arises purely from quantum backreaction; we deal with the case separately.
@@
this term vanishes due the boundary condition on the brane $t^{\prime}(r,0)=0$ and $G(0)=1$. Physically, it can be understood as the quantum corrections due to bulk matter fields Emparan:2021hyr , and in this particular case, it vanishes due to the boundary condition on the brane.
@@
We notice that the leading order quantum correction to the complexity is the same as that of a neutral black hole Emparan:2021hyr , as it does not depends explicitly on $q$. This is justified as explicit charge $q$ only appears at $\mathcal{O}(\ell^{2})$ in (10), we do not expect to see any explicit charge $q$ dependence at $\mathcal{O}(\ell)$. At this order, the only charge dependence can be seen through parameters $\mu$ and $\Delta$.
@@ 4.1 Late time growth of complexity
In this section we focus on the late time regime and study the complexity growth. At late times, $r_{\text{min}}$ is determined by extremizing the conserved momentum in $r$ Barbon:2019tuq ; Barbon:2020olv ; Barbon:2020uux ,
@@
It can be seen from the expressions above that the late time growth of complexity of dyonic black holes are same as that of neutral black holes Emparan:2021hyr up to leading order quantum correction due to $\text{CFT}_{3}$, while the charge $q$ dependence is hidden in $M$ and $\mu$.
@@ 4.2 Quantum dressed defects
In $q\rightarrow 0$, this agrees with the neutral case Emparan:2021hyr . As the charge is $q=\frac{\mu}{2}$ in extremal limit Climent:2024nuj , the complexity growth vanishes. This agrees with the previous observations for a classical 4D Reissner-Nordström black hole in the extremal limit Carmi:2017jqz .
@@ 5.1 Regularized WDW patch
In general, the action over the WDW patch is diverging; there are two standard ways to regularize it: either to put a cutoff surface near the asymptotic boundary, say at $r=r_{\text{max}}$, or consider the WDW to end on the regularized surface itself. In the first regularization picture, we have a GHY contribution from the cutoff surface and two joint contributions from the joint surfaces created from the intersection of the null boundaries of WDW patch and the cutoff surface, while in the second picture, there is no GHY contribution and we get a single joint contribution from the intersection of the future and past null boundaries of WDW patch. Here, we consider the second approach to regularization, in addition to that, rather than considering the full WDW patch till the asymptotic boundary of the $\text{AdS}_{4}$ C-metric at $rx=$-$\ell$, we clip off the WDW at $r=\infty$, denoted by $\widetilde{W}_{\tilde{t}}$. We have shown explicitly in the appendix B that if we consider the full WDW patch, it contributes a constant shift in the total complexity and leaves the late time growth unaffected. To simplify the discussion, we consider $\widetilde{\mathcal{W}}_{\tilde{t}}$ as the relevant object that describes the holographic complexity of the degrees of freedom within the interior of the quBTZ black hole. In particular, we will determine
@@ 5.2 Action computation
Figure 7: Left: The violet points are ‘standard joint surfaces’ in the bulk for a constant $(x,\phi)$ section of the geometry. Right: The violet lines are ‘new joint surfaces’ for a constant $\phi$ section of the brane. The violet lines represent two-dimensional surfaces that extend in the $\phi$ direction along the brane.
@@
where, $n^{(\pm)}_{\mu}$ are normal vectors on the boundaries of the WDW patch in $x>0$ and $x<0$ regions, respectively111111given by, $n_{\mu}=dt\pm\frac{dr}{H(r)}$ for $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ respectively. As the normal vectors $n^{(+)}$ and $n^{(-)}$ are continuous across the brane i.e. $n^{(+)}=n^{(-)}$. This leads to the vanishing contribution to the joint action as the boost vector vanishes $a=0$.
@@ 5.3 Late time growth
and the growth in late time121212we consider the growth in regularized time $\bar{t}=\Delta\tilde{t}$,
@@
It can be noticed that the late time growth of complexity for charged quantum black holes attained by following the CA conjecture agrees with the neutral case Emparan:2021hyr in $q\rightarrow 0$ limit up to a factor [^13] Due to the absence of a GHY contribution from the singularity which gives a non-vanishing contribution in the neutral quantum black hole case.. A significant feature of this result is the prefactor of the mass $M$ term, which is dependent on the parameter $x_{1}$; this distinct it from the previous observations for classical black holes Reynolds:2016rvl ; Brown:2015bva that shows the agreement between CV and CA results up to a prefactor dependent on the spacetime dimension. Another significant point to notice is that the rate of complexity vanishes for a purely magnetic charged case. This point has been previously discussed in Goto:2018iay ; the details of which are given in the section below.
@@ 5.4 Effects of Maxwell boundary term
In Goto:2018iay , it has been shown that for charged black holes, a boundary term for spacelike/timelike boundaries can be added in the action contribution to the complexity.
@@
Also, it can be noticed that for $x_{1}$ large, the known Lloyd bound for dyonic black holes Ovgun:2018jbm is satisfied.
@@ 5.5 Quantum dressed defects
Therefore, the late time growth of complexity in this case gives vanishing results up to $\mathcal{O}(\ell)$, which agrees with the CV as well as previous classical observations Carmi:2017jqz .
@@ 6 Discussion
The leading order piece is that of a classical BTZ black hole, and the first quantum corrected piece is the same as the previously observed neutral and rotating quantum blackholes. This represents the universal nature of quantum corrections to the holographic complexity. It has already been observed in Emparan:2021hyr and Chen:2023tpi that the leading order quantum correction to the late time growth of complexity deviates from the expected behavior $TS_{\text{gen}}$. Furthermore, for the quantum dressed dyonic defects, we found a vanishing late time complexity growth, which is consistent, as from the four-dimensional bulk perspective it represents a four-dimensional extremal Reissner-Nordström black holes Carmi:2017jqz .
@@
This result is closely related to the proposed Lloyd bound for dyonic black holes Ovgun:2018jbm , but with an extended set of thermodynamic quantities. There is a recent work Aguilar-Gutierrez:2023ccv which presents a detailed study of the validity of the Lloyd bound in the braneworld setup of $\text{AdS}_{3}$ with JT gravity localized on the brane. We aim to conduct such a detailed analysis with some generalization added to our framework in the near future.
@@
Our findings show that dyonic quantum black hole geometries encode quantum corrections in entanglement entropy and holographic complexity in a universal manner, consistent with and extending earlier studies Emparan:2021hyr ; Chen:2023tpi . In particular, the behavior we observe suggests a deeper unifying structure that links quantum-corrected spacetime geometry with information-theoretic quantities, thereby strengthening the connection between black hole physics and quantum information.
This work also motivates multiple extensions that we intend to approach in the future. One straightforward extension of our study is to include rotation and U(1) charge, as discussed in Bhattacharya:2025tdn . The complexity of the quantum rotating BTZ black hole has already been explored in Chen:2023tpi . A deeper investigation into entanglement and complexity from an effective field theory perspective also remains an interesting problem Couch:2016exn ; Belin:2021bga ; Jorstad:2023kmq .
Another natural modification involves replacing the bath, which, in our study, has played a subsidiary role, with an alternative gravitating system by incorporating a second brane into the setup. Following this approach will allow us to extend our analysis to wedge holography Mollabashi:2014qfa . Moreover, we can extend to dynamical braneworld scenarios, such as DGP brane constructions Dvali:2000hr . Another interesting feature is the study of the first law of complexity in the case of quantum black hole, which states that the variation of complexity depends only on the end point of the trajectory i.e. target state Bernamonti:2019zyy . These extensions could provide new insights into how entanglement and complexity bounds behave within broader holographic frameworks.
@@ Acknowledgements.
We want to thank Subho R Roy for the various discussions and comments. We would like to express our gratitude to Suvankar Dutta, Nabamita Banerjee, Taniya Mandal, Arpita Mitra, and Arpan Bhattacharyya for their valuable comments on the manuscript. We also appreciate Shinji Hirano for his technical feedback on this work. We sincerely acknowledge the initiative of the $\text{ST}^{4}-2024$ workshop”, the platform where the stage of this work is laid out. G.S.P. wishes to extend thanks to Imtak Jeon and Robert de Mello Koch for their hospitality in Huzhou. G.S.P. also expresses appreciation to Subho R. Roy for organizing the visit to IIT Hyderabad during the final stages of the project. The work of S.P. is supported by the ANRF-SERB research project CRG/2023/001120, “Many Facets of Complexity: From Chaos to Thermalization”.
@@ Appendix A Counter term contribution:
The counter term/LMPS term contribution is necessary to take into account to get rid of the ambiguity by reparameterization of the null generator parameter $\lambda_{n}$ Lehner:2016vdi . The counter term is given by,

### v2

@@ 4 Volume complexity
$\mathcal{C}_{V}(t)=2\int_{r_{\text{min}}}^{\infty}dr\int_{0}^{rx_{1}}dz\frac{\ell^{2}}{(\ell+z)^{3}}f(r,z),$ (48)
@@
We can proceed with the extremization of this volume functional. For the extermination, we can think of the volume functional as a Lagrangian $\mathcal{L}(\dot{t},\dot{r},r)$ which is in general a function of $r(\lambda)$ and $t(\lambda)$ but as in this case it is independent of $t$ we will have a constant of motion,
@@
where, $H_{0}(r)=\frac{r^{2}}{l_{3}^{2}}+\kappa$. If we consider $\kappa=-1$, we get the analog of the classical BTZ black hole result Carmi et al. (2017a) with the classical horizon at $r_{+}=\ell_{3}$ hidding the charged defect. Below, we focus on the quantum correction and late-time growth of complexity in this case. For the $\kappa=1$ case, one can proceed similarly, with the only difference coming from the fact that there is no horizon classically i.e. there is no real root of $H(r)$. The horizon arises purely from quantum backreaction; we deal with the case separately.
@@
this term vanishes due the boundary condition on the brane $t^{\prime}(r,0)=0$ and $G(0)=1$. Physically, it can be understood as the quantum corrections due to bulk matter fields Emparan et al. (2022), and in this particular case, it vanishes due to the boundary condition on the brane.
@@
We notice that the leading order quantum correction to the complexity is the same as that of a neutral black hole Emparan et al. (2022), as it does not depends explicitly on $q$. This is justified as explicit charge $q$ only appears at $\mathcal{O}(\ell^{2})$ in (10), we do not expect to see any explicit charge $q$ dependence at $\mathcal{O}(\ell)$. At this order, the only charge dependence can be seen through parameters $\mu$ and $\Delta$.
@@ 4.1 Late time growth of complexity
In this section we focus on the late time regime and study the complexity growth. At late times, $r_{\text{min}}$ is determined by extremizing the conserved momentum in $r$ Barbón et al. (2020); Barbon et al. (2020, 2021),
@@
It can be seen from the expressions above that the late time growth of complexity of dyonic black holes are same as that of neutral black holes Emparan et al. (2022) up to leading order quantum correction due to $\text{CFT}_{3}$, while the charge $q$ dependence is hidden in $M$ and $\mu$.
@@ 4.2 Quantum dressed defects
In $q\rightarrow 0$, this agrees with the neutral case Emparan et al. (2022). As the charge is $q=\frac{\mu}{2}$ in extremal limit Climent et al. (2024), the complexity growth vanishes. This agrees with the previous observations for a classical 4D Reissner-Nordström black hole in the extremal limit Carmi et al. (2017a).
@@ 5.1 Regularized WDW patch
In general, the action over the WDW patch is diverging; there are two standard ways to regularize it: either to put a cutoff surface near the asymptotic boundary, say at $r=r_{\text{max}}$, or consider the WDW to end on the regularized surface itself Carmi et al. (2017b). In the first regularization picture, we have a GHY contribution from the cutoff surface and two joint contributions from the joint surfaces created from the intersection of the null boundaries of WDW patch and the cutoff surface, while in the second picture, there is no GHY contribution and we get a single joint contribution from the intersection of the future and past null boundaries of WDW patch. The exact equivalence of these two regularization approaches was shown in Akhavan and Omidi (2019) by adding counter terms on these cutoff surfaces (timelike boundaries). Also, it has been shown that the UV divergences can be removed completely by adding counter terms on the null boundaries of the WDW patch. Later in Omidi (2020), the equivalence is shown in the case of a two-sided BTZ black hole truncated by a dynamical timelike ETW brane. Here, we consider the second approach to regularization, in addition to that, rather than considering the full WDW patch till the asymptotic boundary of the $\text{AdS}_{4}$ C-metric at $rx=$-$\ell$, we clip off the WDW at $r=\infty$, denoted by $\widetilde{W}_{\tilde{t}}$. We have shown explicitly in the appendix B that if we consider the full WDW patch, it contributes a constant shift in the total complexity and leaves the late time growth unaffected. To simplify the discussion, we consider $\widetilde{\mathcal{W}}_{\tilde{t}}$ as the relevant object that describes the holographic complexity of the degrees of freedom within the interior of the quBTZ black hole. In particular, we will determine
@@ 5.2 Action computation
Figure 7: Left: The violet points are ‘standard joint surfaces’ in the bulk for a constant $(x,\phi)$ section of the geometry. Right: The sky blue lines are ‘new joint surfaces’ for a constant $\phi$ section of the brane. The sky blue lines represent two-dimensional surfaces that extend in the $\phi$ direction along the brane.
@@
where, $n^{(\pm)}_{\mu}$ are normal vectors on the boundaries of the WDW patch in $x>0$ and $x<0$ regions, respectively151515given by, $n_{\mu}=dt\pm\frac{dr}{H(r)}$ for $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ respectively.. As the normal vectors $n^{(+)}$ and $n^{(-)}$ are continuous across the brane i.e. $n^{(+)}=n^{(-)}$. This leads to the vanishing contribution to the joint action as the boost vector vanishes $a=0$.
@@ 5.3 Late time growth
and the growth in late time161616we consider the growth in regularized time $\bar{t}=\Delta\tilde{t}$.,
@@
It can be noticed that the late time growth of complexity for charged quantum black holes attained by following the CA conjecture agrees with the neutral case Emparan et al. (2022) in $q\rightarrow 0$ limit up to a factor [^17] Due to the absence of a GHY contribution from the singularity which gives a non-vanishing contribution in the neutral quantum black hole case.. A significant feature of this result is the prefactor of the mass $M$ term, which is dependent on the parameter $x_{1}$; this distinct it from the previous observations for classical black holes Reynolds and Ross (2017b); Brown et al. (2016b) that shows the agreement between CV and CA results up to a prefactor dependent on the spacetime dimension. Another significant point to notice is that the rate of complexity vanishes for a purely magnetic charged case. This point has been previously discussed in Goto et al. (2019); the details of which are given in the section below.
@@ 5.4 Effects of Maxwell boundary term
In Goto et al. (2019), it has been shown that for charged black holes, a boundary term for spacelike/timelike boundaries can be added in the action contribution to the complexity.
@@
Also, it can be noticed that for $x_{1}$ large, the known Lloyd bound for dyonic black holes Ovgün and Jusufi (2018) is satisfied.
@@ 5.5 Quantum dressed defects
Therefore, the late time growth of complexity in this case gives vanishing results up to $\mathcal{O}(\ell)$, which agrees with the CV as well as previous classical observations Carmi et al. (2017a).
@@ 6 Discussion
The leading order piece is that of a classical BTZ black hole, and the first quantum corrected piece is the same as the previously observed neutral and rotating quantum blackholes. This represents the universal nature of quantum corrections to the holographic complexity. It has already been observed in Emparan et al. (2022) and Chen et al. (2024) that the leading order quantum correction to the late time growth of complexity deviates from the expected behavior $TS_{\text{gen}}$. Furthermore, for the quantum dressed dyonic defects, we found a vanishing late time complexity growth, which is consistent, as from the four-dimensional bulk perspective it represents a four-dimensional extremal Reissner-Nordström black holes Carmi et al. (2017a).
@@
This result is closely related to the proposed Lloyd bound for dyonic black holes Ovgün and Jusufi (2018), but with an extended set of thermodynamic quantities. There is a recent work Aguilar-Gutierrez et al. (2024b) which presents a detailed study of the validity of the Lloyd bound in the braneworld setup of $\text{AdS}_{3}$ with JT gravity localized on the brane. We aim to conduct such a detailed analysis with some generalization added to our framework in the near future.
@@
Our findings show that dyonic quantum black hole geometries encode quantum corrections in entanglement entropy and holographic complexity in a universal manner, consistent with and extending earlier studies Emparan et al. (2022); Chen et al. (2024). In particular, the behavior we observe suggests a deeper unifying structure that links quantum-corrected spacetime geometry with information-theoretic quantities, thereby strengthening the connection between black hole physics and quantum information.
This work also motivates multiple extensions that we intend to approach in the future. One straightforward extension of our study is to include rotation and U(1) charge, as discussed in Bhattacharya et al. (2025). The complexity of the quantum rotating BTZ black hole has already been explored in Chen et al. (2024). A deeper investigation into entanglement and complexity from an effective field theory perspective also remains an interesting problem Couch et al. (2017); Belin et al. (2022); Jørstad et al. (2023).
Another natural modification involves replacing the bath, which, in our study, has played a subsidiary role, with an alternative gravitating system by incorporating a second brane into the setup. Following this approach will allow us to extend our analysis to wedge holography Mollabashi et al. (2014). Moreover, we can extend to dynamical braneworld scenarios, such as DGP brane constructions Dvali et al. (2000). Another interesting feature is the study of the first law of complexity in the case of quantum black hole, which states that the variation of complexity depends only on the end point of the trajectory i.e. target state Bernamonti et al. (2019). These extensions could provide new insights into how entanglement and complexity bounds behave within broader holographic frameworks.
@@ Acknowledgements.
We want to thank Subho R Roy for the various discussions and comments. We would like to express our gratitude to Suvankar Dutta, Nabamita Banerjee, Taniya Mandal, Arpita Mitra, and Arpan Bhattacharyya for their valuable comments on the manuscript. We also appreciate Shinji Hirano for his technical feedback on this work. We sincerely acknowledge the initiative of the $\text{ST}^{4}-2024$ workshop”, the platform where the stage of this work is laid out. G.S.P. wishes to extend thanks to Imtak Jeon and Robert de Mello Koch for their hospitality in Huzhou. G.S.P. also expresses appreciation to Subho R. Roy for organizing the visit to IIT Hyderabad during the final stages of the project. The work of S.P. is supported by the ANRF-SERB research project CRG/2023/001120, “Many Facets of Complexity: From Chaos to Thermalization”. The work of G.S.P. is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12247103.
@@ Appendix A Counter term contribution:
The counter term/LMPS term contribution is necessary to take into account to get rid of the ambiguity by reparameterization of the null generator parameter $\lambda_{n}$ Lehner et al. (2016). The counter term is given by,